tutor can you help with these 2 problems I have done them several times and come up with a different answer each time and none of them are correct

1) 4-(5-55t)=22t+14


2) 5[2-16(a-16)]=16(5a+2)+1378


Answer by stanbon(75887)

1) 4-(5-55t)=22t+14
4-5+55t = 22t+14
-1+55t = 22t+14
33t = 15
t = 15/33
t 5/11
-----------------------

2) 5[2-16(a-16)]=16(5a+2)+1378
5[2-16a+16^2] = 80a+32 + 1378
5[258-16a] = 80a + 1410
1290-80a = 80a+1410
160a = -120
a = -3/4
=======================
Cheers,
Stan H.